I Vector A has magnitade 50 meters and direction, cousterlockwise from exxt, of 13.9 degroes To I decimal place, what are the values of its componens and its y-componentt 2. A vector has a-component -1.8 and y-component 2.3 What are its magnitude and direction to the nearest tenth of a unit? 3. A baseball is thrown horizontally off a cliff with a speed of 16 m/s. To the nearest tenth of a meter, what is the horizontal distance (x) and how far has it fallen (y) from the face of the cliff after 4.2 seconds? 4. A ball is thrown with a speed of 19 m/s and at an angle of 46 degrees. How high does it go above the cliff to the nearest tenth of a meter?
Answer #1

1)

X component = 50 x cos13.9 = 48.5 m

Y component = 50 x sin13.9 = 12.0 m

2)

Magnitude = sqrt(1.8^2 + 2.3^2) = 2.9

Direction = arctan (2.3/-1.8) = 52.0 degree

Direction is 180 - 52 = 128 degree from +ve x axis

Comment in case any doubt please rate my answer ....

Please post other questions seperately ... .